# FacLens: Transferable Probe for Foreseeing Non-Factuality in Large Language Models

### v3

@@ 5.2 Experimental Results
  • Note: Here FacLens denotes FacLens (last token, middle layer). SAT Probe involves a feature extraction step, where attention weights are extracted for probing, taking 132.860s. Before the training FacLens, we extract hidden question representations, a process that takes only 71.856s.
Unsupervised domain adaptation performs well for cross-LLM FacLens. Given an LLM, we train FacLens on the training data of the corresponding domain and directly test the FacLens on the test data of another domain. The results in the upper part of Figure 6 are unsatisfactory. After unsupervised domain adaptation, the cross-LLM FacLens can work well in the target domain, as depicted in the the lower part of Figure 6. We also discuss the choice of the kernel function in Appendix G, and find that linear kernel performs well, indicating that the NFP features derived by $g_{enc}$ are inherently discriminative. Furthermore, we observe that FacLens demonstrates better transferability between LLMs of similar scales. In future work, we will explore more effective methods to enhance FacLens’s transferability between LLMs of significantly different scales.
Question-aligned strategy is necessary to mini-batch training of cross-LLM FacLens. Figure 7 shows that our question-aligned strategy for mini-batch training significantly enhances the performance of cross-LLM FacLens, demonstrating its efficacy in estimating the true distance between $P_{S}\left(\mathbf{Z}\right)$ and $P_{T}\left(\mathbf{Z}\right)$ within randomly sampled mini-batches. Particularly on the NQ released by Google, which consists of questions from real users and thus covers more diverse questions, the estimation of $P_{S}\left(\mathbf{Z}\right)$ and $P_{T}\left(\mathbf{Z}\right)$ is more likely to be affected by the sampling process. Hence, integrating the question-aligned strategy fosters the training process more on the NQ dataset.
@@
Figure 6: Performance of cross-LLM FacLens w/o and w/ DA. In each heatmap, the element in the $i$-th row and $j$-th column corresponds to the $i$-th source domain and the $j$-th target domain.
Figure 7: Evaluation of the question-aligned mini-bath training strategy. The similar trends appear on other pairs of source and target domains.
5.3 Human Evaluation
We developed and deployed a demo for FacLens, which was utilized for conducting human evaluation to assess its performance in practical use. Details of the demo are introduced in Appendix I. We recruited 22 volunteers, consisting of 11 females and 11 males with bachelor degrees or higher, to use our demo and rate its performance. 3 points indicate that the prediction of FacLens is correct, 2 points indicate that FacLens acknowledges its lack of confidence in the prediction result, and 1 point indicates that the prediction of FacLens is incorrect. We received 680 de-duplicated user queries, with 127 (18.7%) receiving 1 point, 70 (10.3%) receiving 2 points, and 483 (71.0%) receiving 3 points. Overall, these results highlight the effectiveness of FacLens in practice.
@@ 6 Conclusion
In this paper, we find that the hidden representations of users’ input questions contain valuable information for identifying potential non-factual responses (i.e., NFP). We also discover that similar NFP patterns emerge in hidden question representations sourced from different LLMs. These findings support our lightweight and transferable NFP model, FacLens, which offers a more efficient approach to leveraging LLMs’ knowledge awareness for NFP. Extensive experiments show the superiority of FacLens, and we hope this work can inspire future research on improving LLMs’ factuality.
@@ Appendix A Statistics of QA Datasets
During the NFP dataset construction, to reduce the false negative samples, we exclude multiple-choice questions because their LLM-generated responses are likely to mention both correct and incorrect answers. Additionally, we exclude questions where the golden answers are three characters or fewer, as such short strings are likely to appear as substrings within unrelated words. PQ initially contains 14,267 questions. After eliminating duplicates and removing the above special questions, 11,360 unique questions are remained. EQ contains 100K questions. We randomly sample 7,200 questions from EQ, ensuring uniform coverage across all question topics. After eliminating duplicates and removing special questions, we retain 7,159 questions from EQ. The full NQ dataset is 42Gb, so we download a simplified development set of NQ. In this paper, we focus on the case of short answers, so we select questions whose answers are comprised of 30 characters or less. Consequently, our dataset comprises 1,244 questions sourced from NQ. Table 3 shows the ratios of positive and negative samples in each NFP dataset, where a pair of QA dataset and LLM corresponds to an NFP dataset.
Table 3: Positive and negative sample ratios in different NFP datasets (%). A NFP dataset is built based on an LLM and a QA dataset. A positive (non-factual) sample indicates the LLM $m$ cannot provide the queried facts in response to the question $q$, whereas a negative (factual) sample indicates the LLM $m$ can provide the queried facts in response to the question $q$.
    PQ     EQ     NQ
    Pos Neg (factual)     Pos Neg (factual)     Pos Neg (factual)
LLaMA2-7B-Chat     74.9 25.1     70.3 29.7     57.2 42.8
LLaMA3-8B-Instruct     65.5 34.5     61.6 38.4     48.2 51.8
Mistral-7B-Instruct-v0.2     73.0 27.0     68.2 31.8     55.5 44.5
Qwen2-1.5B-Instruct     86.2 13.8     80.1 19.9     75.9 24.1
Figure 8: Illustration of different labeling processes for training FacLens.
Appendix B Illustration of the more efficient labeling process enabled by the transferability of FacLens
@@
In Section 4.2, we have analyzed why the transferability of FacLens can reduce overall development costs by lowering the costs of obtaining labels for FacLens training. Figure 8 provides the illustration, where the gray dashed lines indicate that the corresponding steps are omitted.
Appendix C Perplexity (PPL) on an Input Question
@@
This paper regards PPL as a baseline, we predict $y=1$ if the PPL value exceeds a certain threshold. We extend the calculation of PPL to be conducted in each layer to obtain multiple PPL values for a text, and determine the layer based on the NFP performance on labeled data. Formally, PPL on a question calculated in the $\ell$-th layer is formulated as,
$\centering\text{PPL}=\text{exp}\left(\frac{1}{|q|}\sum_{v_{k}\in q}-\log\left(p_{\ell}\left(v_{k}|v_{<k}\right)\right)\right),p_{\ell}\left(v_{k}|v_{<k}\right)=\text{Softmax}\left(m_{\leq\ell}\left(v_{<k}\right)W_{U}\right)_{v_{k}}\@add@centering$ (7)
@@
where $W_{U}$ is the pre-trained unembedding matrix of the LLM $m$ that converts the hidden representations of tokens into distributions over the vocabulary.
Appendix D Hyper-Parameter Settings
@@
[…]
Appendix E Application Limitations of Certain Baselines
In Table 1, Entity-Popularity and SAT Probe are not suitable for certain datasets. Here, we explain the reasons. Entity-Popularity uses Wikipedia page views to approximate the entity popularity. However, EQ and NQ datasets do not provide the relevant Wikipedia page views, and not every subject entity in the two datasets can be uniquely matched to a Wikidata entity. As a result, Entity-Popularity is unsuitable for EQ and NQ. For the baseline SAT Probe, each question is assumed to contain constraint tokens, and the model extracts LLMs’ attention to the constraint tokens to probe factuality. The authors of SAT Probe have restricted the formats of questions to directly identify the constraint tokens. However, extracting constraint tokens from free-form questions can be challenging. For PQ and EQ, which are template-based, obtaining constraint tokens is relatively straightforward. However, SAT Probe is not suitable for NQ, as questions in NQ come from real users and exhibit diverse structures.
The core of SAT Probe is using an LLM’s attention weights to constraint tokens within a question to reflect the LLM’s factual accuracy. In the original paper, the SAT probe is implemented by a linear layer, optimized by logistic regression. To compare the effectiveness of hidden representations and attention weights, we employ the same MLP structure and CE loss for both SAT Probe and FacLens.
Appendix F Performance of FacLens (avg)
Table 4: Prediction performance of FacLens (avg) (AUC %).
@@ Appendix F Performance of FacLens (avg)
Figure 9: Evaluation of kernels used by MMD loss for training cross-LLM FacLens. The results are derived on the NQ datasets. Similar trends are observed on other QA datasets.
Appendix G Kernel Selection for Computing MMD Loss
@@
Appendix H Cross-LLM FacLens for LLMs of Distinct Hidden Dimensions
Both Qwen2-1.5B-Instruct (Yang et al., 2024a) and Gemma-7B-it (Banks & Warkentin, 2024) have different hidden dimensions compared to LLaMA2-7B-Chat, LLaMA3-8B-Instruct, and Mistral-7B-Instruct-v0.2. The hidden dimension of Qwen2-1.5B-Instruct is 1536, and the hidden dimension of Gemma-7B-it is 3072, while the hidden dimension of LLaMA2-7B-Chat, LLaMA3-8B-Instruct, and Mistral-7B-Instruct-v0.2 is 4096. A FacLens specially trained for the source-domain LLM cannot be directly used for a target-domain LLM whose hidden dimension is distinct from that of the source-domain LLM. Hence we introduce a linear layer to reshape the target-domain hidden question representations to match the dimension of the source domain’s hidden question representations, and still adopt Eq. 6 to conduct domain adaptation. In Figure 10, we can see that although two LLMs have different hidden dimensions, the cross-LLM FacLens can work well.
Figure 10: Evaluation of cross-LLM FacLens on LLMs with different hidden dimensions. The questions are from NQ.
Appendix I Demo & Case Study
@@
User Interface of the Demo. We have implemented a demo of FacLens, whose web user interface is shown in Figure 11. In the demo, a user can choose a specific LLM, and then enters a question in the text box. After submitting the question, FacLens will return whether the LLM knows the factual answer. Then the user can decide whether to call the LLM to generate the response. If the user decides to query the LLM, the demo will provide the response generated by the LLM. According to the prediction of FacLens and the LLM’s response, the user can score the performance of FacLens.
Figure 11: Instructions for using the demo of FacLens.
@@
FacLens in the Demo. Taking LLaMA2-7B-Chat (abbreviated as LLaMA2) as the example, we integrate its NFP datasets, i.e., LLaMA2-PQ, LLaMA2-EQ, and LLaMA2-NQ, to train a FacLens. Specifically, we use instances from LLaMA2-PQ, LLaMA2-EQ for training, and use instances from LLaMA2-NQ for validation. That is because NQ, released by Google, consists of questions posed by real users. We set the learning rate to 1e-3, and determine the training epochs according to the performance on the validation set. On the validation set, we use FacLens to predict the probability $\mathbf{p}\left(y=1|\mathbf{x}\right)$ for each instance. Then we calculate the averaged probability $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{pos}$ based on positive instances in the validation set, as well as the averaged probability $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{neg}$ based on negative instances in the validation set. $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{pos}$ and $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{neg}$ are used as the thresholds for predicting whether the LLM knows the factual answers. Clearly, given a question, FacLens predicts the probability of LLaMA2 generating a non-factual response. If the probability is larger than $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{pos}$, the demo outputs “The LLM does not know the factual answer”; If the probability is smaller than $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{neg}$, the demo outputs “The LLM knows the factual answer”; otherwise, the demo outputs “I am not sure if the LLM knows the factual answer”. Similarly, we train FacLens for LLaMA3-8B-Instruct, Mistral-7B-Instruct-v0.2, and Qwen2-1.5B-Instruct, respectively.
@@
Case Study and Future Work. Taking LLaMA2-7B-Chat as the example, we show some good cases at the end of this paper, based on questions from our human evaluation. These questions cover a range of topics, such as movies, TV series, music, brands, games, and science. We observe that FacLens is more likely to mistakenly predict on questions related to periodic events. The recurrence of similar events could impact FacLens ’s judgments. In other words, LLMs could have limited self-awareness in recognizing the nuances of periodic events. We plan to enhance FacLens’s ability to better handle such questions in future work.
Appendix J Limitations and Future Work
This paper assumes access to the parameters of LLMs, which limits the application of FacLens to API-based black-box LLMs. However, with the growing availability of advanced open-source LLMs, research on white-box methods is becoming increasingly important, making our work meaningful. We also recognize the necessity of designing effective and efficient NFP models for black-box LLMs in our future work.
Good Case User Question: Who was the screenwriter for The Toy? Prediction from FacLens: The LLM does not know the factual answer. Main Response from LLM: The screenwriter for the 1982 film “The Toy” is: Tom Mankiewicz Note: The LLM’s response is incorrect. The screenwriter is not Tom Mankiewicz.
Good Case User Question: What is Taylor Swift’s latest album in 2024? Prediction from FacLens: The LLM does not know the factual answer. Main Response from LLM: Taylor Swift has not released any new albums in 2024. Additionally, she has not announced any upcoming albums or releases. Note: The LLM lacks this real-time knowledge. The LLM’s training cutoff date is likely to be before 2024.
Good Case User Question: Where is ACL 2024 held? Prediction from FacLens: The LLM does not know the factual answer. Main Response from LLM: It is not yet confirmed where ACL 2024 will be held, as the location is typically announced several months to a year in advance. Note: The LLM lacks this real-time knowledge. The LLM’s training cutoff date is likely to be before 2024.
@@
Good Case User Question: Who is the director of TV series The Happy Family? Prediction from FacLens: The LLM does not know the factual answer. Main Response from LLM: I apologize, but I don’t have access to information on the director of a specific TV series called “The Happy Family.” Could you please provide more context or details about the series, such as its release date, network, or any other relevant information? This will help me provide a more accurate answer. Note: The LLM acknowledges that it does not know the factual answer.
@@
Good Case User Question: Who wrote the lyrics for the song I Can Do It With A Broken Heart? Prediction from FacLens: The LLM does not know the factual answer. Main Response from LLM: I apologize, but I don’t have access to the lyrics of the song “I Can Do It With A Broken Heart” as it may be a private or unreleased song. Additionally, I cannot provide lyrics for songs that are not publicly available or have not been released. Note: The LLM acknowledges that it does not know the factual answer.
@@
Good Case User Question: Which TV series does the character Sheldon Lee Cooper come from? Prediction from FacLens: The LLM knows the factual answer. Main Response from LLM: The character Sheldon Lee Cooper is from the popular TV series “The Big Bang Theory.” Note: The LLM’s response is correct.
@@
Good Case User Question: Who founded KFC? Prediction from FacLens: The LLM knows the factual answer. Main Response from LLM: KFC, or Kentucky Fried Chicken, was founded by Colonel Harland Sanders. Note: The LLM’s response is correct.
@@
Good Case User Question: Which is heavier, 1kg iron or 1kg cotton? Prediction from FacLens: The LLM knows the factual answer. Main Response from LLM: Both 1 kg of iron and 1 kg of cotton weigh the same amount, 1 kilogram. Note: The LLM’s response is correct.
Good Case User Question: Is the character controlled by the player in the game “The Legend of Zelda: Breath of the Wild” Zelda? Prediction from FacLens: The LLM knows the factual answer. Main Response from LLM: No, the playable character in The Legend of Zelda: Breath of the Wild is Link, not Zelda. Note: The LLM’s response is correct.
Good Case User Question: Who discovered penicillin? Prediction from FacLens: The LLM knows the factual answer. Main Response from LLM: Penicillin was discovered by Scottish scientist Alexander Fleming in 1928. Note: The LLM’s response is correct.
Good Case User Question: Dogs can eat some grapes and chocolate, right? Prediction from FacLens: The LLM knows the factual answer. Main Response from LLM: No, it’s not safe for dogs to eat grapes or chocolate. Dogs can experience toxicity and potentially life-threatening symptoms if they ingest these substances. Note: The LLM’s response is correct.
Good Case User Question: Is it suitable for families with cats to keep oleander? Prediction from FacLens: The LLM knows the factual answer. Main Response from LLM: No, it is not recommended to keep oleander in a home with cats. Oleander is highly toxic to cats and can cause serious health problems, including cardiac problems, tremors, and even death, if ingested. The sap of the oleander plant can also cause skin irritation and allergic reactions in cats. Note: The LLM’s response is correct.
@@
Bad Case User Question: Which team won the women’s volleyball championship at the 2004 Athens Olympics? Prediction from FacLens: The LLM knows the factual answer. Main Response from LLM: The United States women’s volleyball team won the gold medal at the 2004 Athens Olympics. Note: The LLM’s response is incorrect. The correct answer is the Chinese women’s volleyball team.

### v4

@@ 5.1 Experimental Setup
  • Note: Here FacLens denotes FacLens (last token, middle layer). SAT Probe involves a feature extraction step, where attention weights are extracted for probing, taking 132.860s. Before training FacLens, we extract hidden question representations, a process that takes only 71.856s.
Table 2: Efficiency evaluation of FacLens (seconds).
5.2 Experimental Results
Probing hidden question representations for NFP in fact-seeking QA yields promising results. As reported in Table 1, FacLens exhibits promising performance compared to the baselines, and it favors hidden representations from LLMs’ middle layers. Specifically, Self-Evaluation and LoRA can be regarded as two special representation-based NFP methods. While both perform well, FacLens, with much fewer trainable parameters, offers higher efficiency and yields results comparable to, or even surpassing, those of Self-Evaluation and LoRA. Compared to SAT Probe, FacLens demonstrates that hidden representation could be more useful than attention weights for NFP. Furthermore, we conduct main experiments on HotpotQA Yang et al. (2018), a more complex multi-hop QA dataset, in Appendix D.4. The results show that FacLens continues to perform well. We also developed a demo to collect more questions and provide the case study in Appendix E.
Question-level modeling is more effective than focusing on specific tokens. Inspired by entity-centric studies, we propose FacLens-ent, which feeds the representations of input entity tokens into FacLens. [^6] Entities, e.g., persons, locations, organizations, are identified using the Stanza NLP Package (Qi et al., 2020). As shown in Table 1, FacLens consistently surpasses FacLens-ent, Entity-Popularity, Self-Familiarity, and SAT Probe, suggesting that overemphasizing specific tokens of the input question may mislead the predictions.
@@ 5.2 Experimental Results
Figure 6: Performance of cross-LLM FacLens w/o and w/ DA. In each heatmap, the element in the $i$-th row and $j$-th column represents the performance on the $j$-th target domain, with label information transferred from the $i$-th source domain.
Figure 7: Evaluation of the question-aligned mini-bath training strategy. Similar trends appear on other pairs of source and target domains.
FacLens stands out for its efficiency. Taking the LLaMA2-PQ NFP dataset as an example, which includes 2,272 questions for training, 1,136 questions for validation, and 7,952 questions for testing, Table 2 reports the average training time per epoch and the average prediction time per question for each method. We can see that fine-tuning an LLM (i.e., Self-Evaluation and LoRA) incurs significant computational costs. Importantly, if the LLM has been fine-tuned for a specific task, its ability on other tasks can be compromised (Yang et al., 2024b). As for prediction, FacLens runs much faster than Self-Familiarity because Self-Familiarity involves multiple conversations with the LLM. As FacLens uses the middle-layer hidden question representations, it runs faster than Prompting, PPL, LoRA, and Self-Evaluation that involve more layers in the LLM. SAT Probe extracts attention weights across all layers and attention heads, so FacLens runs faster during prediction.
Unsupervised domain adaptation performs well for cross-LLM FacLens. Given an LLM, we train FacLens with the training data of the corresponding domain and directly test it on the test data of another domain. The results in the upper part of Figure 6 are unsatisfactory. After unsupervised DA (MMD loss with linear kernel), the cross-LLM FacLens can work much better in the target domain, as depicted in the the lower part of Figure 6. Furthermore, we observe that FacLens shows better transferability between LLMs of similar scales. In future work, we will explore more effective methods to enhance FacLens’s transferability between LLMs of very different scales.
Question-aligned strategy is necessary to mini-batch training of cross-LLM FacLens. Figure 7 shows that our question-aligned strategy for mini-batch training significantly enhances the performance of cross-LLM FacLens. Particularly on the NQ dataset released by Google, which consists of questions from real users and covers more diverse questions, the estimation of $P_{S}\left(\mathbf{Z}\right)$ and $P_{T}\left(\mathbf{Z}\right)$ is more likely to be influenced by the sampling process in a mini-batch. Hence, integrating the question-aligned strategy fosters the training process more on NQ.
@@ 6 Conclusion
In this paper, we find that the hidden representation of a fact-seeking question contains valuable information for identifying potential non-factual responses (i.e., NFP). We also discover that similar NFP patterns emerge in hidden question representations sourced from different LLMs. These findings support our lightweight and transferable NFP model, FacLens, which enables more efficient development and application. We hope this work can inspire future research on LLMs’ factuality.
Limitations
This paper assumes access to the parameters of LLMs, which limits the application of FacLens to API-based black-box LLMs. With the growing availability of advanced open-source LLMs, research on white-box methods is becoming increasingly important, making our work meaningful. However, we also recognize the necessity of designing effective and efficient NFP models for black-box LLMs in our future work.
@@ Appendix A Statistics of QA Datasets
During the NFP dataset construction, to reduce the false negative samples, we exclude multiple-choice questions because their LLM-generated responses are likely to mention both correct and incorrect answers. Moreover, we exclude questions where the golden answers are three characters or fewer, as such short strings are likely to appear as substrings within unrelated words. PQ initially contains 14,267 questions. After eliminating duplicates and removing the above special questions, 11,360 unique questions remained. EQ contains 100K questions. We randomly sample 7,200 questions from EQ, ensuring uniform coverage across all question topics. After eliminating duplicates and removing special questions, we retain 7,159 questions from EQ. The full NQ dataset is huge (about 42GB), so we download a simplified development set of NQ. In this paper, we focus on the case of short answers, so we select questions whose answers are comprised of 30 characters or less. Hence, we include 1,244 questions sourced from NQ. Table 3 shows the ratios of positive and negative samples in each NFP dataset, where a pair of QA dataset and LLM corresponds to an NFP dataset.
Appendix B More Efficient Process of Training Data Construction
@@
In Section 4.2, we analyzed why the transferability of FacLens can reduce overall development costs for multiple LLMs by lowering the costs of obtaining labels for FacLens training. Figure 8 provides the illustration, where the gray dashed lines indicate that the corresponding steps are omitted.
Appendix C Experimental Settings
C.1 Perplexity (PPL) on a Question
@@
We regard the PPL of a fact-seeking question as a baseline. In specific, we predict $y=1$ if the PPL value exceeds a certain threshold. We extend the calculation of PPL to be conducted in each layer to obtain multiple PPL values for a text and determine the layer based on the NFP performance on labeled data. Formally, PPL on a question calculated in the $\ell$-th layer is formulated as,
$\text{PPL}=\exp\left(\frac{1}{|q|}\sum_{v_{k}\in q}-\log\left(p_{\ell}\left(v_{k}|v_{<k}\right)\right)\right)$ (7)
$p_{\ell}\left(v_{k}|v_{<k}\right)=\text{Softmax}\left(m_{\leq\ell}\left(v_{<k}\right)W_{U}\right)_{v_{k}}$ (8)
@@
where $q$ is a fact-seeking question, $v_{k}$ is the $k$-th token in $q$, $v_{<k}$ represents the set of tokens preceding the $k$-th token, and $W_{U}$ is the pre-trained unembedding matrix of the LLM $m$ that converts the hidden token representations into distributions over the vocabulary.
C.2 Hyper-Parameter Settings
Our experiments are conducted based on 4 * 80G NVIDIA Tesla A800 GPUs. We implement the encoder $g_{enc}$ of FacLens by a 3-layer MLP, setting the dimension of each MLP layer to 256. We use the Adam optimizer with weight decay 1e-4. The hyper-parameters determined on the validation set include: the training epochs (set the maximum epochs to 100), and the learning rate $\in$ {1e-3, 1e-4} for single-LLM FacLens. Considering that the training questions from NQ are relatively small, we set the learning rate of FacLens to 1e-4 on NFP datasets derived from NQ. The default learning rate of cross-LLM FacLens is set to 1e-5. Due to the memory limitation, we minimize the MMD loss via mini-batch training with a batch size of 64.
@@ C.2 Hyper-Parameter Settings
Figure 9: Performance comparison of FacLens (ante-hoc NFP) with INSIDE and SAPLMA (post-hoc NFD) on EQ and NQ.
Figure 10: Performance comparison between $f_{m}$ and $f_{mix}$ on EQ and NQ. Similar performance suggests no significant concept shift across different domains.
Figure 11: Visualizations of NFP features.
C.3 Why Certain Baselines Are Unsuitable for the QA Datasets
In Table 1, Entity-Popularity and SAT Probe are not suitable for certain datasets. Here, we explain the reasons. Entity-Popularity uses Wikipedia page views to approximate the entity’s popularity. However, EQ and NQ datasets do not provide relevant Wikipedia page views, and not every subject entity in the two datasets can be uniquely matched to a Wikidata entity. As a result, Entity-Popularity is unsuitable for EQ and NQ. For the baseline SAT Probe, each question is assumed to contain constraint tokens, and the model extracts LLMs’ attention to the constraint tokens to probe factuality. The authors of SAT Probe have restricted the formats of questions to directly identify the constraint tokens. However, extracting constraint tokens from free-form questions can be challenging. For PQ and EQ, which are template-based, obtaining constraint tokens is relatively straightforward. However, SAT Probe is not suitable for NQ, as questions in NQ come from real users and exhibit diverse structures.
The core of the SAT Probe is using an LLM’s attention weights to constraint tokens within a question to reflect the LLM’s factual accuracy. In the original paper, the SAT probe is implemented by a linear layer, optimized by logistic regression. To compare the effectiveness of hidden representations and attention weights, we employ the same MLP structure and CE loss for both SAT Probe and FacLens.
Appendix D Supplementary Experiments
D.1 Ante-Hoc NFP vs. Post-Hoc NFD
In Figure 9, we supplement the performance comparison between hidden representation-based ante-hoc and post-hoc methods on EQ and NQ. The results further highlight that FacLens (ante-hoc NFP method) has the potential to outperform the post-hoc NFD methods.
@@ Appendix D Supplementary Experiments
Table 4: Prediction performance of FacLens (avg) (AUC %).
  LLaMA2  LLaMA3  Mistral  Qwen2
PPL    55.2    55.2    54.4    53.5
Prompting    62.5    61.0    63.0    62.1
Entity-Popularity    –    –    –
SAT Probe    –    –    –
Self-Familiarity    55.3    56.8    54.5    53.7
LoRA (Parameter-Efficient FT)    72.9    68.0    66.9    70.8
Self-Evaluation (Fully FT)    75.0    69.3    71.1    72.1
FacLens (last token, last layer)    74.3    68.6  74.1    72.1
[…]
FacLens (last token, middle layer)    75.5  66.9    74.9  71.4
Table 5: Prediction performance of different NFP methods on HotpotQA (AUC %).
D.2 Demonstrations of FacLens’s Transferability
In Section 4.2, we compare FacLens $f_{m}$, trained on an individual domain, with FacLens $f_{mix}$, trained on the mixture domain. The results show that $f_{mix}$ exhibits comparable performance to $f_{m}$ on the test set of the corresponding individual domain. We supplement the results based on questions from EQ and NQ in Figure 10, further demonstrating no significant concept shifts between domains.
Besides, we supplement more visualizations of the NFP features in Figure 11, further demonstrating that a unified classification boundary can be applied to different LLMs in the NFP task.
D.3 Performance of FacLens (avg)
Figure 12: Evaluation of cross-LLM FacLens on LLMs of different hidden dimensions. Questions are from NQ.
D.4 Evaluation on HotpotQA
Table 5 shows the prediction performance of different NFP methods on HotpotQA. The observations are consistent with that in Table 1, indicating FacLens can also handle the more complex fact-seeking questions.
D.5 Cross-LLM FacLens for LLMs of Distinct Hidden Dimensions
Both Qwen2-1.5B-Instruct (Yang et al., 2024a) and Gemma-7B-it (Banks and Warkentin, 2024) have different hidden dimensions compared to LLaMA2-7B-Chat, LLaMA3-8B-Instruct, and Mistral-7B-Instruct-v0.2. The hidden dimension of Qwen2-1.5B-Instruct is 1536, and the hidden dimension of Gemma-7B-it is 3072, while the hidden dimension of LLaMA2-7B-Chat, LLaMA3-8B-Instruct, and Mistral-7B-Instruct-v0.2 is 4096. A FacLens specially trained for the source-domain LLM cannot be directly used for a target-domain LLM whose hidden dimension is distinct from that of the source-domain LLM. Hence we introduce a linear layer to reshape the target-domain hidden question representations to match the dimension of the source domain’s hidden question representations and still adopt Eq. 6 to conduct domain adaptation. In Figure 12, we can see that although two LLMs have different hidden dimensions, the cross-LLM FacLens can work well.
Figure 13: Evaluation of kernels used by MMD loss for training cross-LLM FacLens. The results are derived from the NQ dataset. Similar trends are observed on other QA datasets.
D.6 Kernel Selection for MMD Loss
@@
Appendix E Demo & Case Study
@@
Human Evaluation via Demo. We have implemented a demo of FacLens. In the demo, a user can choose a specific LLM, and then enter a fact-seeking question in the text box. After submitting the question, FacLens will return whether the LLM knows the factual answer. Then the user can decide whether to call the LLM to generate the response. If the user decides to query the LLM, the demo will provide the response generated by the LLM. According to the prediction of FacLens and the LLM’s response, the user can score the performance of FacLens. We recruited 22 volunteers, consisting of 11 females and 11 males with bachelor’s degrees or higher, to use our demo and rate its performance. 3 points indicate that the prediction of FacLens is correct, 2 points indicate that FacLens acknowledges its lack of confidence in the prediction result, and 1 point indicates that the prediction of FacLens is incorrect. We received 680 de-duplicated user queries, with 127 (18.7%) receiving 1 point, 70 (10.3%) receiving 2 points, and 483 (71.0%) receiving 3 points.
@@
FacLens in the Demo. Taking LLaMA2-7B-Chat (abbreviated as LLaMA2) as the example, we integrate its NFP datasets, i.e., LLaMA2-PQ, LLaMA2-EQ, and LLaMA2-NQ, to train a FacLens. Specifically, we use instances from LLaMA2-PQ, and LLaMA2-EQ for training, and instances from LLaMA2-NQ for validation. That is because NQ, released by Google, consists of questions posed by real users. We set the learning rate to 1e-3, and determine the training epochs based on the performance on the validation set. On the validation set, we use FacLens to predict the probability $\mathbf{p}\left(y=1|\mathbf{x}\right)$ for each instance. Then we calculate the averaged probability $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{pos}$ based on positive instances in the validation set, as well as the averaged probability $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{neg}$ based on negative instances in the validation set. $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{pos}$ and $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{neg}$ are used as the thresholds for predicting whether the LLM knows the factual answers. Clearly, given a question, FacLens predicts the probability of LLaMA2 generating a non-factual response. If the probability is larger than $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{pos}$, the demo outputs “The LLM does not know the factual answer”; If the probability is smaller than $\mathbf{\overline{p}}\left(y=1|\mathbf{x}\right)_{neg}$, the demo outputs “The LLM knows the factual answer”; otherwise, the demo outputs “I am not sure if the LLM knows the factual answer”. Similarly, we train FacLens for LLaMA3-8B-Instruct, Mistral-7B-Instruct-v0.2, and Qwen2-1.5B-Instruct, respectively.
@@
Case Study and Future Work. Taking LLaMA2-7B-Chat as an example, we show some good cases at the end of this paper, based on questions from our human evaluation. These questions cover a range of topics, such as movies, TV series, music, brands, games, and science. We observe that FacLens works well on most questions, and it is more likely to mistakenly predict on questions related to periodic events. The recurrence of similar events could impact FacLens ’s judgments. In other words, LLMs could have limited self-awareness in recognizing the nuances of periodic events. We plan to enhance FacLens’s ability to better handle such questions in future work.
Good Case User Question: Who was the screenwriter for The Toy?
Prediction from FacLens: The LLM does not know the factual answer.
Main Response from LLM: The screenwriter for the 1982 film “The Toy” is: Tom Mankiewicz
Note: The LLM’s response is incorrect. The screenwriter is not Tom Mankiewicz.
Good Case User Question: Who is the director of TV series The Happy Family?
Prediction from FacLens: The LLM does not know the factual answer.
@@
Main Response from LLM: I apologize, but I don’t have access to information on the director of a specific TV series called “The Happy Family.” Could you please provide more context or details about the series, such as its release date, network, or any other relevant information? This will help me provide a more accurate answer.
Note: The LLM acknowledges that it does not know the factual answer.
Good Case User Question: Who wrote the lyrics for the song I Can Do It With A Broken Heart?
Prediction from FacLens: The LLM does not know the factual answer.
@@
Main Response from LLM: I apologize, but I don’t have access to the lyrics of the song “I Can Do It With A Broken Heart” as it may be a private or unreleased song. Additionally, I cannot provide lyrics for songs that are not publicly available or have not been released.
Note: The LLM acknowledges that it does not know the factual answer.
@@
Good Case User Question: Which TV series does the character Sheldon Lee Cooper come from?
Prediction from FacLens: The LLM knows the factual answer.
Main Response from LLM: The character Sheldon Lee Cooper is from the popular TV series “The Big Bang Theory.”
Note: The LLM’s response is correct.
Good Case User Question: Who founded KFC?
Prediction from FacLens: The LLM knows the factual answer.
@@
Main Response from LLM: KFC, or Kentucky Fried Chicken, was founded by Colonel Harland Sanders.
Note: The LLM’s response is correct.
Good Case User Question: Which is heavier, 1kg iron or 1kg cotton?
Prediction from FacLens: The LLM knows the factual answer.
@@
Main Response from LLM: Both 1 kg of iron and 1 kg of cotton weigh the same amount, 1 kilogram.
Note: The LLM’s response is correct.
Bad Case User Question: Which team won the women’s volleyball championship at the 2004 Athens Olympics?
Prediction from FacLens: The LLM knows the factual answer.
@@
Main Response from LLM: The United States women’s volleyball team won the gold medal at the 2004 Athens Olympics.
Note: The LLM’s response is incorrect. The correct answer is the Chinese women’s volleyball team.
Bad Case User Question: Who won the Turing Award in 2016?
Prediction from FacLens: The LLM knows the factual answer.